Circular No. 7317

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gc IN ANONYMOUS GALAXY J. Maza, University of Chile; and M. Hamuy, University of Arizona,

report the discovery by R. Antezana, on a 20-min unfiltered T-Max 400 film taken by L. Gonzalez on Nov. 10.25 UT with the University of Chile’s Maksutov telescope, of a supernova (m_{pv} ∼ 17.5) located at α = 1^{h}05^{m}10^{s}.87, δ = −20^{o}17′48′′.1 (equinox 2000.0), which is 6′′.7 west and 10′′.7 north of the nucleus of the host galaxy. The supernova was confirmed by L.-G. Strolger (University of Michigan) from B, V, R, I images taken with the Cerro Tololo 1.5-m telescope on Nov. 16. Hamuy obtained a spectrum (range 350–900 nm; resolution 0.3 nm) of SN 1999gc on Nov. 19.13 using the New Technology Telescope at the European Southern Observatory, revealing that the supernova is a type-Ia event, ∼ 3 weeks past V maximum light. Weak emission lines from the host galaxy indicate a redshift z = 0.054. Additional B, V, I images taken with the Cerro Tololo 0.91-m telescope by Antezana and Hamuy on Nov. 24.23 yield B = 19.81 and V = 18.65 (± 0.05).

HD 209458 G. Gonzalez, University of Washington, writes that he obtained differ-

ential CCD photometry of HD 209458 during Nov. 22.07–22.35 UT, using the the Manastash Ridge Observatory 0.76-m telescope, with C. Laws and M. Braunstein. An 8-nm-side filter centered at 673 nm was used to obtain images every ∼ 3 min. They detect a dimming of HD 209458 starting at Nov. 22.28 ± 0.015 day, but they did not observe long enough to see last contact, due to the low altitude of the star. They estimate that the dimming amounted to ∼ 0.028 mag (using two comparison stars that were each ∼ 4 mag fainter than HD 209458).

COMET C/1999 S3 (LINEAR)

Ephemeris extension to IAUC 7285 (elements MPC 36653):

1999/2000 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 28  20^{h}34.^{m}26 +53^{o}45.^{′}3 1.575 1.908 93.^{o}4 31.^{o}1 12.8
Dec. 8  20 27.09 +52 55.6 1.685 1.926 88.2 30.8 13.0
18  20 25.02 +52 45.6 1.788 1.950 84.1 30.1 13.2
28  20 26.38 +53 15.0 1.882 1.981 80.9 29.4 13.3
Jan. 7  20 30.17 +54 22.3 1.964 2.018 78.8 28.5 13.5
17  20 35.67 +56 06.3 2.036 2.061 77.6 27.8 13.7
1999 November 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green